A chemist was preparing a solution that should have included 35 milligrams of a chemical. If she actually used 36.4 milligrams, what was her percentage error ( to the nearest .001%)?
Answer by Alan3354(69443)

A chemist was preparing a solution that should have included 35 milligrams of a chemical. If she actually used 36.4 milligrams, what was her percentage error ( to the nearest .001%)?
--------
The error was 1.4 mg
1.4/35 = 1/25 = 4/100
= 4%